OpenStudy (anonymous):

how do you do this problem ? f+4/g=6 solve for f

jimthompson5910 (jim_thompson5910):

First you multiply both sides by 'g'. Then you subtract 4 from both sides to fully isolate f. \[\Large \frac{f+4}{g} = 6\] \[\Large f+4 = 6g\] \[\Large f = 6g-4\]